# P3H-25-058, TTP25-027

### v1

@@ 2.3 Addressing the $B$ anomalies
As a phase rotation of complex fields, this symmetry also emerges in the gauge sector. Taking $S_{3}$ as an example, although the SM gauge symmetry allows it to couple to both $(\overline{Q}_{L}L_{L}^{c})$ and $(\overline{Q_{L}^{c}}Q_{L})$, the minimal Yukawa sector does not contain the $\overline{Q_{L}^{c}}S_{3}Q_{L}$ term because it is not invariant under $U(1)_{\text{PQ}}$. To suppress these diquark couplings also at the loop level, the PQ symmetry should also be well preserved in the scalar potential. The $\eta_{2}(126_{H})^{4}$ and $\gamma_{2}(45_{H})^{2}(126_{H})^{2}$ terms Bertolini:2012im explicitly break the PQ symmetry. To ensure that the proton decay amplitude is suppressed by $M_{\text{GUT}}^{2}$, the magnitudes of $\gamma_{2}$ and $\eta_{2}$ can be much larger than about $\mathcal{O}\left(\frac{m_{S}}{M_{\text{GUT}}}\right)^{2}$, where the $m_{S}$ denotes lightest mass among $S_{3},\overline{S}_{1}$ and $\widetilde{S}_{1}$ [^3] Since the hierarchy could be large, we omit the possible loop factors, which can relax the bound by $(16\pi^{2})^{n}$, where $n$ is the loop order.. If $m_{S}\sim$ TeV, $\gamma_{2}$ and $\eta_{2}$ are constrained extremely small. By itself, this does not directly introduce a new hierarchy puzzle, since $\gamma_{2}=\eta_{2}=0$ enhances the PQ symmetry. However, the $H_{u}-H_{d}$ mixing term is proportional to $\gamma_{2}M_{\text{GUT}}^{2}$. To achieve a sizable mixing angle $\beta$, the mass of $H_{d}$ cannot be far larger than $\sqrt{\gamma_{2}}M_{\text{GUT}}\lesssim\mathcal{O}(m_{\text{S}})$. This requires additional fine-tuning, while it enriches the low-energy spectrum with an extra Higgs doublet. Its charged component can contribute to $b\rightarrow c\tau\nu$ Crivellin:2012ye ; Crivellin:2013wna ; Iguro:2022uzz ; Blanke:2022pjy and further relax the tension between explaining $R(D^{(*)})$ and avoiding $B\rightarrow K\overline{\nu}\nu$ constrains.
To conclude, $126_{H}$ contains six types of LQs: $S_{3},S_{1},\overline{S}_{1},\widetilde{S}_{1}$ and $R_{2},\widetilde{R}_{2}$. Among them, $S_{1}$ is excluded from being light due to the proton decay constrains. $R_{2}$ is disfavored by both Yukawa and gauge coupling unification. $\overline{S}_{1}$ only couples to $\nu_{R}$. On the other hand, $S_{3},\widetilde{S}_{1}$ and $\widetilde{R}_{2}$ can address the $B$ anomalies, and a consistent explanation typically requires a combination of two or all of them.
@@ 3.1 $b-\tau$ masses
Here, no inhomogeneous terms arise because the $U(1)_{\text{PQ}}$ symmetry ensures each LQ only couples to a single type of fermion bilinear. This feature further simplifies the RG equation compared to the general case Fedele:2023rxb . The beta function for every coupling is always proportional to itself. Moreover, the running receives negative contributions from the gauge interactions and positive contributions from Yukawa couplings.
@@
Figure 2: RG evolution of the third-generation charged fermions masses from $10^{2}$ to $10^{16}$ GeV. The solid (dotted) lines indicate the scenario with (without) LQs at TeV scale. We choose $y_{b}=y_{t}=-\frac{1}{3}y_{\tau}=\frac{1}{2\sqrt{3}}y_{1}=\frac{1}{2\sqrt{3}}y_{2}=\frac{1}{2\sqrt{3}}y_{3}=0.56$ at GUT scale, and $\tan\beta=42$. The gray vertical lines indicate $M_{\text{GUT}}$ and the light LQ threshold.
We illustrate how light LQs from $126_{H}$ change the mass running in Figure 2. We assume exact Yukawa unification at $M_{\text{GUT}}$ (Eq (14) holds exactly) and fix the free inputs $y_{t}(M_{\text{GUT}})=0.56,~\tan\beta=42$. The solid lines indicate how the $b$-, $\tau$-, $t$- masses evolve from $10^{16}$ GeV down to $10^{2}$ GeV. For $1~\text{TeV}<\mu<10^{16}~\text{GeV}$, we use the running equation shown in Eq (15). Below $1~\text{TeV}$, we directly interpolate the SM running data included in the SMDR package. Martin:2019lqd . The resulting $b$-, $\tau$-, $t$-masses at 1 TeV can agree reasonably well with the values extracted from the SM. While the matching at 1 TeV would be perfect if the solid lines were continuous, the gaps visible in the plots are not problematic. At leading-log level, TeV-scale LQs can reduce the $b-\tau$ mass tension from an $\mathcal{O}(1)$ discrepancy to a suppressed $\mathcal{O}(\epsilon)$ mismatch. For comparison, we also show the evolution without TeV-scale LQs, by plotting the SM running parameters from Ref. Martin:2019lqd with dots for $\mu>1$ TeV. The red and blue dotted curves still diverge significantly at $10^{16}$ GeV, indicating the relation $m_{b}=m_{\tau}/3$ cannot be reached if the theory below $M_{\text{GUT}}$ is SM alone.
We also explore scenarios with only $S_{3}+\widetilde{R}_{2}$ and only $S_{3}+\widetilde{S}_{1}$. By choosing different values of $y_{t}(M_{\text{GUT}})$ and $\tan\beta$, both scenarios yield successful Yukawa unification at $10^{16}$ GeV and reasonably good matching relationships at 1 TeV. Related plots are shown in Appendix A. This universal feature suggests that the improved $b-\tau$ unification is not an accidental outcome of a particular LQ choice, but rather a model-independent effect from colored scalar fields. We understand the underlying reason as follows. If the IR theory is simply the SM, $(m_{\tau}/3)$ is too small and remains nearly a constant. $m_{b}$ decreases as the energy scale $\mu$ goes up, mainly due to QCD effects, but not fast enough to meet $(m_{\tau}/3)$ at $M_{\text{GUT}}$. As shown in Eq (15), the additional homogeneous terms in the $\beta$-function of $y_{\tau}$ are always positive and rapidly drive $y_{\tau}$ towards a Landau-pole at UV. Although a similar behavior is also seen for $y_{b}$, its growth is slower because $y_{\tau}$ receives an extra enhancement by a factor of $N_{c}=3$.
@@ 3.2 LQ-fermion couplings
The LQ-fermion couplings $y_{1},y_{2},y_{3}$ are also outcomes from the RG evolution. They are correlated at $M_{\text{GUT}}$ by Eq (14) and evolve together with $y_{t},y_{b},y_{\tau}$ according to Eq (15). Since $y_{t}(M_{\text{GUT}})$ is fixed by the measured value of $m_{t}$, $y_{1},y_{2}$ and $y_{3}$ in the range $1~\text{TeV}<\mu<M_{\text{GUT}}$ become predictions of the minimal GUT, as shown in the left panel of Figure 3. A fixed-point behavior Pendleton:1980as ; Hill:1980sq emerges: all LQ-fermion couplings drop rapidly near $M_{\text{GUT}}$ and remain nearly constant for $\mu\lesssim 10^{5}\sim 10^{10}$ GeV. This behavior is similar to what recently found in Ref. Fedele:2023rxb , that the LQ-fermions couplings approach IR fixed-points of $0.5\sim 1.0$. Interestingly, the LQ-type considered here is different, suggesting that this feature is a rather general consequence from RG evolution.
Eq (15) contains no inhomogeneous terms, so the IR fixed points can be solved analytically. Neglecting $g_{1},g_{2}$ for simplicity and assuming $g_{3}$ varies sufficiently slowly, we find the fixed point solution:
@@
As a result, LQ-fermion couplings are numerically close to $g_{3}$ and $y_{t}$ at TeV-scale. Together with Ref. Fedele:2023rxb , this supports the scalar LQ explanation of the $B$ anomalies, because $\mathcal{O}(1)$ couplings are typically required to account for the data. Moreover, we find $y_{2}$ and $y_{3}$ almost coincide at all scales. Although this is an accidental result, it can be understood by setting $y_{2}=y_{3}$ and neglecting $g_{1},g_{2},y_{t},y_{b}$ in Eq (15):
@@ 3.3 Emerging flavor mixing angles
A new tension arises. The $b\rightarrow c,s$ transition generally requires sizable non-diagonal Yukawa couplings. However, Eq (4) contains only one Yukawa matrix $Y_{126}$, and one can always choose a diagonal basis. Consequently, no flavor mixing can arise at any scale. Where does the necessary flavor mixing come from? One may construct extended models that introduce sizable off-diagonal LQ Yukawa couplings while leaving the other sectors unaffected or only mildly modified. Such an approach, nevertheless, is ad hoc and sacrifices the elegance of minimality. We reflect on the robustness of the prediction and realize that although the absolute flavor conserving solution remains stable under the RG equations following the SM Ma:1979cw ; Sasaki:1986jv ; Grzadkowski:1987tf , it becomes unstable when light LQs are introduced. If the flavor-conserving structure at $M_{\text{GUT}}$ is not exact, the RG flow from UV to IR can amplify deviations and generate sizable flavor violating interactions at low energies. In this sense, flavor mixing may not mainly originate from explicit next-to-minimal model building, but rather manifest as an emergent phenomenon. The UV theory remains elegantly simple, while complexity arise dynamically through evolving towards IR. This approach is widely applied to address the large neutrino mixing angles Babu:1993qv ; Tanimoto:1995bf ; Balaji:2000gd ; Balaji:2000ma ; Hagedorn:2004ba .
@@
Strictly speaking, the flavor violating interactions cannot be exactly zero at $M_{\text{GUT}}$. Unlike gauge coupling unification, the diagonal flavor structure does not directly come from the SO(10) symmetry, but rather from the assumption of minimality. Potential corrections may arise due to additional scalars that couple to the fermions via loops Witten:1979nr ; Bajc:2004hr ; Bajc:2005aq , or high-dimensional operators Ellis:1979fg generated by for instance gravity or by vector-like Fermions. Without loosing generality, we do not specify the underlying source but add the following flavor-violating interactions to Eq (13):
Here, $Q_{L}^{2}=(c_{L},s_{L})$ donates the second-generation left-handed quark doublet. The flavor basis are chosen such that the $S_{3}$ and $\widetilde{R}_{2}$ couplings are aligned to $Q_{L}^{3}$ and $b_{R}$ at $M_{\text{GUT}}$, respectively. The $t_{R}-c_{R}$ rotation remains unphysical, so the $\overline{Q_{L}^{3}}c_{R}H_{u}$ term is not included. Furthermore, mixings involving leptons or first-generation fermions are omitted, as they are irrelevant to the processes under consideration. $\mathcal{L}_{Y}^{\epsilon}$ is defined at $\Lambda\gtrsim M_{\text{GUT}}$, where $\Lambda$ is the scale that the new dynamics breaking the minimal GUT prediction arises.
Below $M_{\text{GUT}}$, the RG equations for the flavor violating coupling $(\epsilon^{bs}y_{b})$ is given by:
@@
Here, we set $\epsilon_{1}^{bs}=0$ for simplicity, and neglect higher order flavor violating terms. Comparing with $y_{b}$ running in Eq (15), the $\left(\frac{y_{1}^{2}}{2}+y_{2}^{2}\right)$ term is absent in Eq (21). This is because the LQ couplings are third-generation specific and do not contribute to the self-energy corrections for $s_{R}$, as shown in Figure 4. While evolving from UV to IR, $(\epsilon^{bs}y_{b})$ decrease more slowly than $y_{b}$, and the mixing parameters $\epsilon^{bs}$ effectively increase. This behavior is further supported by the negative terms arising in its explicit running equation:
@@
If $\epsilon^{bs}$ is zero at a given scale, it remains exactly zero for all $\mu$ since the right-hand side of Eq (22) vanishes. However, any slight deviation $\epsilon^{bs}$ from zero would be driven towards sizable values as $\mu$ decrease. As a consequence, the $s_{R}-b_{R}$ mixing is enhanced at IR relative to their UV quantities by a factor of $\left(\frac{M_{\text{GUT}}}{\text{TeV}}\right)^{\frac{y_{1}^{2}+2y_{2}^{2}}{32\pi^{2}}}$. This leads to flavor misalignment between the Higgs and LQ sector. Upon choosing the physical basis in which the $H_{d}$ Yukawa coupling is diagonal, the enhanced flavor mixings are then transferred to the $\widetilde{S}_{1}$ and $\widetilde{R}_{2}$ couplings. [^5] There are additional contributions to the flavor misalignment. The flavor violating LQ interaction itself, defined as zero in Eq (20), also receives radiative correction. The beta-function contains an inhomogeneous term proportional to $\epsilon^{bs}y_{b}^{2}$. It is much smaller than the $\epsilon^{bs}\left(\frac{y_{1}^{2}}{2}+y_{2}^{2}\right)$ contribution, because Eq (14) tells that $y_{b}^{2}$ is $12$ times smaller than $y_{1}^{2}$ or $y_{2}^{2}$ at $M_{\text{GUT}}$. For simplicity, we neglect this sub-leading effect.
@@
Figure 5: RG evolution of $\epsilon^{bs}$ with $y_{t}(M_{\text{GUT}})=0.58$. The dashed line lies in the region that the Yukawa couplings become non-perturbative and may not reflect the physical reality.
@@
We illustrate our idea in Figure 5. Keeping $y_{t}(M_{\text{GUT}})=0.58$ as initial condition, the solid line shows the evolution of $\epsilon^{bs}$. Flavor conserving is clearly not a stable solution to the equations, because $\epsilon^{bs}$ increase as $\mu$ decrease. The shortage is that the evolving below $M_{\text{GUT}}$ is too slow: achieving $\epsilon^{bs}\approx 0.5$ at TeV typically requires its GUT scale value be at least $0.35$, which does not fit as a perturbation. However, we find $\epsilon^{bs}$ changes sufficiently fast above $M_{\text{GUT}}$. One reason is that the $(6_{c},1_{L},1_{R})$, $(\overline{10_{c}},1_{L},3_{R}),$ and $(15_{c},2_{L},2_{R})$ multiples contained in $126_{H}$ all contribute to the $b_{R}$ self-energy above the threshold. We include their effects according to the matching condition and RG equations from Ref Aulakh:2002zr ; Meloni:2016rnt ; Aydemir:2018cbb . Another reason is the absolute Yukawa coupling strength becomes large. Using the standard one-loop running equations for gauge coupling $g_{10}$ and Yukawa coupling $Y_{126}$ Vaughn:1978st ; Vaughn:1981qi , we find $Y_{126}$ increases rapidly and approaches the non-perturbative regime when $\mu\gtrsim 10^{17}$ GeV. [^6] Large $Y_{126}$ induces a sizable $g_{10}$ through two-loop effects, while a sizable $g_{10}$ in turn slows down the running of $Y_{126}$ or can even drive it to decrease. If the scalar self-couplings are excluded, asymptotic freedom can be restored in deep UV Vaughn:1981qi . When $\epsilon^{bs}$ lies in the non-perturbative region, we show it with dashed lines, indicating it is only for illustration and may not reflect the physical reality. Therefore, if tiny $\epsilon^{bs}$ originate from physics at scale $\Lambda\gtrsim 10M_{\text{GUT}}$, the running effects yield a sizable value at IR.

### v2

@@ 2.3 Addressing the $B$ anomalies
As a phase rotation of complex fields, this symmetry also emerges in the gauge sector. Taking $S_{3}$ as an example, although the SM gauge symmetry allows it to couple to both $(\overline{Q}_{L}L_{L}^{c})$ and $(\overline{Q_{L}^{c}}Q_{L})$, the minimal Yukawa sector does not contain the $\overline{Q_{L}^{c}}S_{3}Q_{L}$ term because it is not invariant under $U(1)_{\text{PQ}}$. To suppress these diquark couplings also at the loop level, the PQ symmetry should also be well preserved in the scalar potential. The $\eta_{2}(126_{H})^{4}$ and $\gamma_{2}(45_{H})^{2}(126_{H})^{2}$ terms Bertolini:2012im explicitly break the PQ symmetry. To ensure that the proton decay amplitude is suppressed by $M_{\text{GUT}}^{2}$, the magnitudes of $\gamma_{2}$ and $\eta_{2}$ can be much larger than about $\mathcal{O}\left(\frac{m_{S}}{M_{\text{GUT}}}\right)^{2}$, where the $m_{S}$ denotes lightest mass among $S_{3},\overline{S}_{1}$ and $\widetilde{S}_{1}$. [^1] Since the hierarchy could be large, we omit the possible loop factors, which can relax the bound by $(16\pi^{2})^{n}$, where $n$ is the loop order. If $m_{S}\sim$ TeV, $\gamma_{2}$ and $\eta_{2}$ are constrained extremely small. By itself, this does not directly introduce a new hierarchy puzzle, since $\gamma_{2}=\eta_{2}=0$ enhances the PQ symmetry. However, the $H_{u}-H_{d}$ mixing term is proportional to $\gamma_{2}M_{\text{GUT}}^{2}$. To achieve a sizable mixing angle $\beta$, the mass of $H_{d}$ cannot be far larger than $\sqrt{\gamma_{2}}M_{\text{GUT}}\lesssim\mathcal{O}(m_{\text{S}})$. This requires additional fine-tuning, while it enriches the low-energy spectrum with an extra Higgs doublet. Its charged component can contribute to $b\rightarrow c\tau\nu$ Crivellin:2012ye ; Crivellin:2013wna ; Iguro:2022uzz ; Blanke:2022pjy and further relax the tension between explaining $R(D^{(*)})$ and avoiding $B\rightarrow K\overline{\nu}\nu$ constrains.
To conclude, $126_{H}$ contains six types of LQs: $S_{3},S_{1},\overline{S}_{1},\widetilde{S}_{1}$ and $R_{2},\widetilde{R}_{2}$. Among them, $S_{1}$ is excluded from being light due to the proton decay constrains. $R_{2}$ is disfavored by both Yukawa and gauge coupling unification. $\overline{S}_{1}$ only couples to $\nu_{R}$. On the other hand, $S_{3},\widetilde{S}_{1}$ and $\widetilde{R}_{2}$ can address the $B$ anomalies, and a consistent explanation typically requires a combination of two or all of them. It is worth to remark here that the precise flavor data is not strongly correlated to Yukawa unification. The coefficients for the effective operators addressing the B anomalies are suppressed by $1/M_{\text{LQ}}^{2}$. Here, $M_{\text{LQ}}$ represents the mass of the relevant LQ. Varying $M_{\text{LQ}}$ around the TeV scale (to fit the B anomalies) only brings sub-leading noise to the $b-\tau$ mass relationship, because the effects of RG evolution is proportional to $\log(M_{\text{LQ}}/M_{\text{GUT}})$. Therefore, we do not examine in detail the regions preferred by the precise flavor data, but only stress the B anomalies are data-driven motivations for light LQs.
@@ 3.1 $b-\tau$ masses
Here, $g_{1},g_{2},$ and $g_{3}$ are the coupling strengths of $U(1)_{Y},SU(2)_{L},$ and $SU(3)_{c}$, respectively. No inhomogeneous terms arise because the $U(1)_{\text{PQ}}$ symmetry ensures each LQ only couples to a single type of fermion bilinear. This feature further simplifies the RG equation compared to the general case Fedele:2023rxb . The beta function for every coupling is always proportional to itself. Moreover, the running receives negative contributions from the gauge interactions and positive contributions from Yukawa couplings.
@@
Figure 2: RG evolution of the third-generation charged fermions masses from $10^{2}$ to $10^{16}$ GeV. The solid (dotted) lines indicate the scenario with (without) LQs at TeV scale. We fix $y_{t}(M_{\text{GUT}})=0.56$ to get the correct value of $m_{t}$ and Yukawa unification implies $\tan\beta=42$. The gray vertical lines indicate $M_{\text{GUT}}$ and the light LQ threshold.
We illustrate how light LQs from $126_{H}$ change the mass running in Figure 2. We assume exact Yukawa unification at $M_{\text{GUT}}$ (Eq. (14) holds exactly) and fix the two input parameters to $y_{t}(M_{\text{GUT}})=0.56$ and $\tan\beta=42$. The solid lines indicate how the $b,\tau,t$ masses evolve from $10^{16}$ GeV down to $10^{2}$ GeV. For $1~\text{TeV}<\mu<10^{16}~\text{GeV}$, we use the running equation shown in Eq. (15). Below $1~\text{TeV}$, we directly interpolate the SM running data included in the SMDR package. Martin:2019lqd . The resulting $b,\tau,t$ masses at 1 TeV can agree reasonably well with the values extracted from the SM. While the matching at 1 TeV would be perfect if the solid lines were continuous, the gaps visible in the plots are not problematic. At leading-log level, TeV-scale LQs can reduce the $b-\tau$ mass tension from an $\mathcal{O}(1)$ discrepancy to a suppressed $\mathcal{O}(\epsilon)$ mismatch. For comparison, we also show the evolution without TeV-scale LQs, by plotting the SM running parameters from Ref. Martin:2019lqd with dots for $\mu>1$ TeV. The red and blue dotted curves still diverge significantly at $10^{16}$ GeV, indicating the relation $m_{b}=m_{\tau}/3$ cannot be reached if the theory below $M_{\text{GUT}}$ is SM alone.
We also explore scenarios with only $S_{3}+\widetilde{R}_{2}$ and only $S_{3}+\widetilde{S}_{1}$. For both scenarios we are able to find values for $y_{t}(M_{\text{GUT}})$ and $\tan\beta$ leading to successful Yukawa unification at $10^{16}$ GeV and reasonably good matching relationships at 1 TeV. Related plots are shown in Appendix A. This universal feature suggests that the improved $b-\tau$ unification is not an accidental outcome of a particular LQ choice, but rather a model-independent effect from colored scalar fields. We understand the underlying reason as follows. If the IR theory is simply the SM, $(m_{\tau}/3)$ is too small and remains nearly a constant. $m_{b}$ decreases as the energy scale $\mu$ goes up, mainly due to QCD effects, but not fast enough to meet $(m_{\tau}/3)$ at $M_{\text{GUT}}$. As shown in Eq. (15), the additional homogeneous terms in the $\beta$-function of $y_{\tau}$ are always positive and rapidly drive $y_{\tau}$ towards a Landau-pole at UV. Although a similar behavior is also seen for $y_{b}$, its growth is slower because $y_{\tau}$ receives an extra enhancement by a factor of $N_{c}=3$.
@@ 3.2 LQ-fermion couplings
The LQ-fermion couplings $y_{1},y_{2},y_{3}$ are also outcomes from the RG evolution. They are correlated at $M_{\text{GUT}}$ by Eq. (14) and evolve together with $y_{t},y_{b},y_{\tau}$ according to Eq. (15). Since $y_{t}(M_{\text{GUT}})$ is fixed by the measured value of $m_{t}$, $y_{1},y_{2}$ and $y_{3}$ in the range $1~\text{TeV}<\mu<M_{\text{GUT}}$ become predictions of the minimal GUT, as shown in the left panel of Figure 3. A fixed-point behavior Pendleton:1980as ; Hill:1980sq emerges: all LQ-fermion couplings drop rapidly near $M_{\text{GUT}}$ and remain nearly constant for $\mu\lesssim 10^{5}\sim 10^{10}$ GeV. This behavior is similar to what recently found in Ref. Fedele:2023rxb , that the LQ-fermions couplings approach IR fixed-points of $0.5\sim 1.0$. Interestingly, the LQ-type considered here is different, suggesting that this feature is a rather general consequence from RG evolution.
Eq. (15) contains no inhomogeneous terms, so the IR fixed points can be solved analytically. Neglecting $g_{1},g_{2}$ for simplicity and assuming $g_{3}$ varies sufficiently slowly, we find the fixed point solution:
@@
As a result, LQ-fermion couplings are numerically close to $g_{3}$ and $y_{t}$ at TeV-scale. Together with Ref. Fedele:2023rxb , this supports the scalar LQ explanation of the $B$ anomalies, because $\mathcal{O}(1)$ couplings are typically required to account for the data. Moreover, we find $y_{2}$ and $y_{3}$ almost coincide at all scales. Although this is an accidental result, it can be understood by setting $y_{2}=y_{3}$ and neglecting $g_{1},g_{2},y_{t},y_{b}$ in Eq. (15):
@@ 3.3 Emerging flavor mixing angles
The implementation of flavor mixing still needs to be discussed. Although the $b\rightarrow c,s$ transitions generally require sizable non-diagonal Yukawa couplings, Eq. (4) contains only one Yukawa matrix $Y_{126}$, and one can always choose a flavor diagonal basis. Consequently, no flavor mixing can arise at any scale in the minimal set-up. Thus where does the necessary flavor mixing come from? One may construct extended models that introduce sizable off-diagonal LQ Yukawa couplings while leaving the other sectors unaffected or only mildly modified. Such an approach, nevertheless, is ad hoc and sacrifices the elegance of minimality.
From a different perspective, it is also necessary to include flavor mixing for a complete analysis. One physical reason is that the flavor-conserving limit is not protected by gauge symmetry, but comes out as a result of minimality, i.e. because the theory contains only one unique Yukawa coupling matrix. Since $M_{\text{GUT}}$ is close to the Planck scale, new particles related to gravity are believed to lie not far beyond $M_{\text{GUT}}$ with additional Yukawa couplings. Potential flavor violating corrections can arise from additional scalars that couple to the fermions via loops Witten:1979nr ; Bajc:2004hr ; Bajc:2005aq , or higher-dimensional operators Ellis:1979fg generated by vector-like fermions or other gravity related GUT multiplets. A second reason is our new finding in this work that although the flavor-conserving limit is stable under the SM RGEs Ma:1979cw ; Sasaki:1986jv ; Grzadkowski:1987tf , it becomes unstable when light LQs are introduced. If flavor conservation is not exact at $M_{\text{GUT}}$, the RG flow can amplify deviations and generate larger flavor violating interactions at low energies. [^3] A similar example is the amplification of the neutrino mixing angles in the MSSM Babu:1993qv ; Tanimoto:1995bf ; Balaji:2000gd ; Balaji:2000ma ; Hagedorn:2004ba by the RGEs. The flavor mixing can manifest itself as an emergent phenomenon with a small seed at deep UV, while complexity arises dynamically through the evolution towards the IR.
Due to the two reasons mentioned above, we must include the most general and relevant flavor violating effects in a model independent way. The minimal GUT predicts the flavor conserving interaction Eq. (13) only — an over-restrictive result. Despite this, no matter which theory lies behind it, all possible 2nd-to-3rd generation quark flavor violating effects can be captured by the following four new interacting terms:
Here, $Q_{L}^{2}=(c_{L},s_{L})$ denotes the second-generation left-handed quark doublet. The flavor basis is chosen such that the $S_{3}$ and $\widetilde{R}_{2}$ couplings are aligned to $Q_{L}^{3}$ and $b_{R}$, respectively, which can be done without loss of generality. Once $H_{u},H_{d}$ acquire VEVs the diagonalisation of the quark mass matrix provides the needed flavor-changing couplings for $S_{3}$ and $\widetilde{R}_{2}$. Since the $t_{R}-c_{R}$ rotation remains unphysical, we do not need to include the $\overline{Q_{L}^{3}}c_{R}H_{u}$ term. Mixings involving leptons or first-generation fermions are omitted, as they are irrelevant for the $b\to s$ and $b\to c$ processes under consideration. In the quark sector with only 2nd and 3rd generations, Eq. (20) serves as the maximally possible flavor extension, as long as no other light scalar particles are included. (Our approach to add the maximally allowed flavor-breaking term $\mathcal{L}_{Y}^{\epsilon}$ is a well-established methodology, in analogy with the soft SUSY-breaking Lagrangian for the MSSM which is also chosen maximal to cover all possible SUSY-breaking scenarios.) Therefore, we do not specify the underlying mechanism of flavor breaking, but instead explore Eq. (20) since it is generic and covers all possible models.
We believe it is important to remark here, that adding general flavor mixing terms in Eq. (3.9) does not render the GUT scale relation $m_{\tau}=3m_{b}$ arbitrary. This relation is predicted by the $\text{diag}(1,1,1,-3)$ vacuum structure of the Higgs fields in the adjoint representation of $SU(4)_{c}$, which is the relevant symmetry to unify quarks and leptons. $m_{\tau}=3m_{b}$ originates from the Higgs field with $SU(4)_{c}\times SU(2)_{L}\times SU(2)_{R}$ quantum numbers $(15,2,2)$. To change $m_{\tau}=3m_{b}$ at the GUT scale, one needs a $(1,2,2)$ Higgs field or an effective operator containing $(15,1,1)\times(15,2,2)$. By contrast, any extensions to the minimal Yukawa sector leads to flavor violation. For instance, two $(15,2,2)$ Higgs fields (which can originate from two $126_{H}$) incorporate flavor but still predict $m_{\tau}/m_{b}=3$. Furthermore, we will argue that the RG evolution which is responsible for the enhancement of flavor mixing does not change $m_{\tau}/m_{b}$ because of $SU(4)_{c}$ symmetry. Contrary to the case of flavor mixing, large corrections to $m_{\tau}/m_{b}$ can only originate from the mass splitting of particles inside the $SU(4)_{c}$ multiplets, which is realized with TeV-scale leptoquarks in our work, or a non-minimal SO(10) Yukawa sector with large couplings, while small perturbations at the GUT scale cannot change this.
@@
Since Eq. (20) contains the general flavor mixing effects, renormalization does not lead to new physical counterterms, ensuring that Eq. (13) and Eq. (20) are well defined at all scales. We then study how the flavor violating corrections evolve from $\Lambda$, the scale at which the new dynamics correcting the minimal GUT ansatz arises (typically $\Lambda\gtrsim M_{\text{GUT}}$), toward the TeV scale where the light LQs lie. Taking $\epsilon^{bs}$ as an example, the RG equations from $M_{\text{GUT}}$ to TeV scale are given by:
@@
Here, we set $\epsilon_{1}^{bs}=0$ for simplicity, and neglect higher order flavor-violating terms. Comparing with the $y_{b}$ running in Eq. (15), the $\left(\frac{y_{1}^{2}}{2}+y_{2}^{2}\right)$ term is absent in Eq. (21). This is because the LQ couplings are third-generation specific and do not contribute to the self-energy corrections for $s_{R}$, as shown in Figure 4. During the evolution from UV to IR, $(\epsilon^{bs}y_{b})$ decreases more slowly than $y_{b}$, and the mixing parameter $\epsilon^{bs}$ effectively increases. This behavior is further enhanced by the negative terms arising in its explicit RG equation:
@@
If $\epsilon^{bs}$ is zero at a given scale, it remains exactly zero for all scales $\mu$ since the right-hand side of Eq. (22) vanishes. However, any slight deviation of $\epsilon^{bs}$ from zero would be driven towards sizable values as $\mu$ decreases. As a consequence, the $s_{R}-b_{R}$ mixing is enhanced in the IR relative to the UV value by a factor of $\left(\frac{M_{\text{GUT}}}{\text{TeV}}\right)^{\frac{y_{1}^{2}+2y_{2}^{2}}{32\pi^{2}}}$. This leads to a flavor misalignment between the Higgs-Yukawa and LQ sectors. Upon choosing the physical basis in which the $H_{d}$ Yukawa coupling is diagonal, the enhanced flavor mixings are then transferred to the $\widetilde{S}_{1}$ and $\widetilde{R}_{2}$ couplings. [^4] There are additional contributions to the flavor misalignment. The flavor-violating LQ interaction itself, chosen zero in Eq. (20), also receives radiative correction. The beta-function contains an inhomogeneous term proportional to $\epsilon^{bs}y_{b}^{2}$. It is much smaller than the $\epsilon^{bs}\left(\frac{y_{1}^{2}}{2}+y_{2}^{2}\right)$ contribution, because Eq. (14) tells that $y_{b}^{2}$ is $12$ times smaller than $y_{1}^{2}$ or $y_{2}^{2}$ at $M_{\text{GUT}}$. For simplicity, we neglect this sub-leading effect. The physical bottom quark mass also receives a correction which is quadratic in $\epsilon^{bs}$ from the diagonalization of the down-type quark mass matrix. Similar corrections also apply to $m_{\tau}$. Both together may further refine the $b-\tau$ unification, which is welcome to fix e.g. the small mismatches in Fig. 2 and the $S_{3},\widetilde{S}_{1}$ scenario of Fig. 6.
@@
Figure 5: RG evolution of $\epsilon^{bs}$ with $y_{t}(M_{\text{GUT}})=0.58$. The dashed line lies in the region where the Yukawa couplings become non-perturbative and thus may not reflect the physical reality.
@@
We illustrate our idea in Figure 5. Keeping $y_{t}(M_{\text{GUT}})=0.58$ as the initial condition, the solid line shows the evolution of $\epsilon^{bs}$. The preservation of flavors is clearly not a stable solution of the RG equations, because $\epsilon^{bs}$ increases as $\mu$ decreases. The drawback is that the evolution below $M_{\text{GUT}}$ is too slow: Achieving $\epsilon^{bs}\approx 0.5$ at the TeV scale typically requires its GUT scale value be at least $0.35$, which one may consider to be too large for a perturbation. However, we find $\epsilon^{bs}$ changes sufficiently fast above $M_{\text{GUT}}$. One reason is that the $(6_{c},1_{L},1_{R})$, $(\overline{10_{c}},1_{L},3_{R}),$ and $(15_{c},2_{L},2_{R})$ multiplets contained in $126_{H}$ all contribute to the $b_{R}$ self-energy above the threshold. We include their effects according to the matching condition and RG equations from Ref. Aulakh:2002zr ; Meloni:2016rnt ; Aydemir:2018cbb . Another reason is that the absolute value of the Yukawa coupling becomes large. Using the one-loop running equations for gauge and Yukawa couplings above $M_{\text{GUT}}$ Vaughn:1978st ; Vaughn:1981qi ; Fang:2025hlc , we find $Y_{126}$ increases rapidly and approaches the non-perturbative regime when $\mu\gtrsim 10^{17}$ GeV. [^5] Large $Y_{126}$ induces a sizable $g_{10}$ through two-loop effects, while a sizable $g_{10}$ in turn slows down the running of $Y_{126}$ or can even drive it to decrease. If the scalar self-couplings are excluded, asymptotic freedom can be restored in the deep UV Vaughn:1981qi . When $\epsilon^{bs}$ lies in the non-perturbative region, we show it with a dashed line and emphasize that this is only for illustration and may not reflect the physical reality. It is worth noticing that this enhancement originates from the self-energy corrections to a single $16_{F}$, which are therefore universal to all 3rd generation quarks and leptons. The RG evolution from about $10M_{\text{GUT}}$ to $M_{\text{GUT}}$ cannot amplify any perturbation to $m_{\tau}/m_{b}=3$, due to the protection by $SU(4)_{c}$ symmetry. In conclusion, if physics at the scale $\Lambda\gtrsim 10\,M_{\text{GUT}}$ gives a tiny correction to our minimal theory, the RG effects yield sizable flavor mixings in the IR in general, while TeV-scale LQ are still needed to achieve realistic $b-\tau$ mass relationship.